Operations Report – January 9th



Crew 272 Operations Report 9-1-2023

SOL: 8

Name of person filing report: Mason Kuhn

Non-nominal systems: Wiring in EVA Suit 7

Notes on non-nominal systems: I have located an issue with the wiring after opening the back of the suit. It could be a damaged wire or a semi-functional soldered connection. I will send a separate email detailing my findings because I have a few attachments I would like Mission Support to look through.


Spirit rover used: No

Hours: 206.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 111.5

Beginning charge: 100%

Ending charge: 67%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 219.0

Beginning charge: 100%

Ending charge: 64%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: Opportunity and Curiosity were used on today’s EVA. Nominal performance from both.

Summary of Hab operations: Per Mission Support’s request during 272’s mission training, I have replaced the zip-tie connections on the 2 piece suits (from the backpack to the helmet collar) with the provided steel cable ties. I tried on the suit to test its comfortability, and it was similar to slightly more comfortable (as well as more durable). I also opened the backs of suits 7 (non-nominal) and 9 (nominal) to compare the two electrical wiring systems. An issue was located in suit 7’s wiring.

WATER USE: 22.9 gallons

Water (static tank): 226.5 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits 1, 5, 9, and 10 were used during the EVA to Kissing Camel Ridge. All suits performed nominally with airflow and minimal fogging. All suits have been cleaned and are currently charging. There were no issues with radios or headsets on the EVA, and the new headsets are an improvement! All radios that were used have been charged and removed from the charging dock.

Summary of GreenHab operations: Monitored and watered GreenHab crops and experimental crops. The lighting system was found to be unplugged, and it was plugged back in last night by mission support. The lights did turn on last night and are fully functional.

WATER USE: 7 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 6g beans, 30g dill

Summary of Science Dome operations: No crew activity, there was Mission Support activity.

Dual split: On (from approximately 8:00 pm to 8:00 am)

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support:

  • Please see my next email for information about suit 7

Best Regards,

Mason Kuhn

Crew 272 Engineer

Copyright © The Mars Society. All rights reserved. | Main Site